which of the following acids should be used to prepare a buffer with ph 9.3?

A.HbrO Ka= 2.3x10^-9
B.CH3CO2H ka=1.8x10^-5
C. HCO2H Ka= 1.7x10^-4
D. HCN ka= 4.9x10^-10
Answer #1

As we know Henderson- Hasselbalch som of PH = pka + log([A-] THAI where [A]: concentratiem conjugate base & [Hal- concentratiHCOAH, ka 17x10-4 : pka= -log (1.78164) =-{ log (1.7) + log 104] - (0.23-4) 1pka = 3.77) @ HCN, ka =4.9/10-10 . pka = -log (4